OpenStudy (anonymous):

If the area of the smaller triangle is 28 cm2, what is the area of the larger triangle? 168 cm2 112 cm2 84 cm2 56 cm2

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

so can being like multiplie of 28 for example 2*28 =56

OpenStudy (anonymous):

Can you please help me?

OpenStudy (anonymous):

with other questions

jhonyy9 (jhonyy9):

or 56+28=84

OpenStudy (anonymous):

it was 112

jhonyy9 (jhonyy9):

84+28=112

OpenStudy (anonymous):

Would you be able to help me with other questions please?

jhonyy9 (jhonyy9):

or 112+56=168

jhonyy9 (jhonyy9):

so yes but you need write it on open question column

OpenStudy (anonymous):

and tag u in it ?

jhonyy9 (jhonyy9):

ok